6(x+1) = 12(x-3)
then you get
6x+6=12x-36
My question at this point is: How do you know which side to start with. Not sure whether to start with +6 or -36 or 6x or 12x.

Answer by longjonsilver(2297)

it doesn't matter...
either collect x terms together first:
-6x+6=-36
-6x = -42
x=(-42)/(-6)
x=7
or collect the pure numbers together first:
6x = 12x-42
-6x = -42
x = (-42)/(-6)
x = 7
jon